OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x-5y=-31 -12x+15y=93

OpenStudy (anonymous):

-3(4x-5y)=93 -12x-12x+15y-15y=93-93 -24x=0 x=0 -12x+15y+12y+15x=93+93 30y=186 186/30=y